# A Unified Approach to Routing and Cascading for LLMs

### v1

@@ Abstract
The widespread applicability of large language models (LLMs) has increased the availability of many fine-tuned models of various sizes targeting specific tasks. Given a set of such specialized models, to maximize overall performance, it is important to figure out the optimal strategy for selecting the right model for a given user query. An effective strategy could drastically increase overall performance and even offer improvements over a single large monolithic model. Existing approaches typically fall into two categories: routing, where a single model is selected for each query, and cascading, which runs a sequence of increasingly larger models until a satisfactory answer is obtained. However, both have notable limitations: routing commits to an initial model without flexibility, while cascading requires executing every model in sequence, which can be inefficient. Additionally, the conditions under which these strategies are provably optimal remain unclear. In this work, we derive optimal strategies for both routing and cascading. Building on this analysis, we propose a novel approach called cascade routing, which combines the adaptability of routing with the cost-efficiency of cascading. Our experiments demonstrate that cascade routing consistently outperforms both routing and cascading across a variety of settings, improving both output quality and lowering computational cost, thus offering a unified and efficient solution to the model selection problem. [^1] Code available at https://github.com/eth-sri/cascade-routing.
@@ 1 Introduction
Two primary strategies have been proposed to solve model selection. The first, routing, directs each input query to a specific model from a set of available models (Chen et al., 2022; Liu et al., 2024), as illustrated in Fig. 1(a). This approach is particularly effective when different expert LLMs are needed for different tasks, enabling the selection of the most suitable expert for each query. The second strategy, cascading, processes an input query through a sequence of increasingly larger models, stopping when a model produces an answer deemed sufficiently good (Chen et al., 2023; Varshney and Baral, 2022), as illustrated in Fig. 1(b). Cascading is particularly valuable for handling queries of varying difficulty, as it allows simpler queries to be addressed by smaller, less expensive models while reserving more complex queries for larger models.
Restrictive Conditions
Despite their utility, both routing and cascading impose significant restrictions on the model selection process. In routing, the initial selection of a model is final, preventing any reconsideration after the initial decision. In cascading, each query must sequentially pass through all models in the chain until a suitable answer is found, with no option to reroute to a potentially better model. Therefore, a less restrictive strategy that combines the strengths of both routing and cascading could offer significant performance improvements.
Lack of Theoretical Understanding
Furthermore, the conditions under which current routing and cascading strategies are optimal, are not well understood. For routing, an extensive proof is required just to show that current strategies are close to optimal (Chen et al., 2022), while the theoretical analysis of current cascading algorithms is based on restrictive assumptions and does not provide optimality guarantees (Chen et al., 2023; Varshney and Baral, 2022). This lack of theoretical understanding hinders the development of more effective model selection strategies.
@@
To address these limitations, we first derive optimal routing and cascading strategies by framing them as linear optimization problems aimed at maximizing output quality while remaining within a given cost budget. For routing, this optimal strategy is close to the one obtained by prior work, while for cascading we derive a new strategy that is provably better than existing approaches. Building on this theoretical analysis, we propose a new paradigm called cascade routing, which generalizes both routing and cascading. As illustrated in Fig. 1(c), cascade routing initially routes a query to any available model but keeps rerouting to different models until a model produces an answer of sufficient quality. We prove the optimality of our cascade routing strategy and show that it offers significantly more flexibility in processing a query.
@@ Results
We evaluate cascade routing on a wide range of tasks, demonstrating that it significantly outperforms both routing and cascading. Notably, cascade routing consistently improves performance by $1\%$ to $4\%$ across all settings on the popular RouterBench benchmark (Hu et al., 2024), which represents a relative improvement over a naive baseline by an additional $13\%$ to $80\%$. Furthermore, we show that our new cascading strategy outperforms existing cascading strategies by up to $2\%$, validating that our theoretical analysis leads to practical improvements over prior work.
Contributions
Our main contributions are as follows:
@@
  • We conduct a thorough evaluation, demonstrating that cascade routing consistently outperforms the baselines (Section 5).
@@ 2 Routing as Linear Optimization
We derive an optimal routing strategy to select the best model for a given query, providing detailed proofs for all statements in this section in App. A. We will use the analysis presented here to develop the optimal cascading and cascade routing strategies in Section 3 and Section 4, respectively.
Routing Strategy
@@
A routing strategy selects a model by sampling from the distribution $s(x)$ for each query $x$. In prior work, routing strategies were restricted to be deterministic, i.e., $s_{i}(x)\in\{0,1\}$ (Chen et al., 2022; Hu et al., 2024). In contrast, we propose using a more general probabilistic routing strategy that enables a better solution and an easier theoretical analysis.
@@
However, since these functions are unknown in practice, we need estimators $\hat{q}_{i}(x)$ and $\hat{c}_{i}(x)$ that approximate the output quality and cost of querying model $m_{i}$ on input $x$. Estimators for $q_{i}(x)$ can be created using small classifiers trained to predict model accuracy, as done in prior work (Hu et al., 2024; Shnitzer et al., 2023). $\hat{c}_{i}(x)$ can be estimated by tokenizing the input query and determining the average output length of the model on a query. Then, we can use API-specific costs per token to estimate the cost of running a model on a query.
@@
The optimal routing strategy $s_{\textsc{opt}}$ for a given cost budget $B$ is the solution to the optimization problem
$\max_{s}$ $\mathbb{E}_{x\in\mathcal{X},i\in\{1,\dots,k\}}(s_{i}(x)\hat{q}_{i}(x))$ (1)
s.t. $\mathbb{E}_{x\in\mathcal{X},i\in\{1,\dots,k\}}(s_{i}(x)\hat{c}_{i}(x))\leqslant B.$
We now explain how to solve this linear optimization problem. Intuitively, the optimal routing strategy optimizes the cost-quality tradeoff $\tau_{i}(x,\lambda)=\hat{q}_{i}(x)-\lambda\hat{c}_{i}(x)$, where $\lambda\in\mathbb{R}^{+}$ controls the balance between quality and cost based on the budget $B$. For each query $x$, the model that achieves the highest value of $\tau_{i}(x,\lambda)$ is selected.
More formally, for a given $\lambda\in\mathbb{R}^{+}$, we define two deterministic routing strategies: $s_{\textsc{min}}^{\lambda}(x)$, which selects the cheapest model achieving the optimal cost-quality tradeoff, and $s_{\textsc{max}}^{\lambda}(x)$, which selects the most expensive model achieving this tradeoff. The optimal routing strategy $s_{\textsc{opt}}$ is then determined by the following theorem:
@@
For a given cost budget $B$, there exists a $\lambda\in\mathbb{R}^{+}$ and a $\gamma\in[0,1]$ such that the optimal routing strategy $s_{\textsc{opt}}$ equals $\gamma s_{\textsc{min}}^{\lambda}+(1-\gamma)s_{\textsc{max}}^{\lambda}$. Furthermore, all routing strategies that have an expected cost that is exactly equal to $B$ and can be written as a convex combination of $s_{\textsc{min}}^{\lambda^{\prime}}$ and $s_{\textsc{max}}^{\lambda^{\prime}}$ for some $\lambda^{\prime}\in\mathbb{R}^{+}$ achieve the same optimal quality.
Since $\gamma$ is often not equal to 00 or $1$, the optimal routing strategy is not deterministic and instead selects a model probabilistically. Therefore, prior work that only considered deterministic routing strategies (Chen et al., 2022; Hu et al., 2024) cannot express the routing strategy from Theorem 1 and fall back to the near-optimal $s_{\textsc{min}}^{\lambda}$ instead.
Due to the second part of Theorem 1, we only need to find a set of hyperparameters $\lambda$ and $\gamma$ that achieve the desired cost budget. To determine these parameters, we estimate the cost of each strategy using a validation dataset $D$ that is representative of the query distribution $\mathcal{X}$. Since the cost of $s_{\textsc{min}}^{\lambda}$ increases as $\lambda$ decreases (see App. A), we use a binary search to find the appropriate value of $\lambda$. $\gamma$ is determined by interpolating between the costs of $s_{\textsc{min}}^{\lambda}$ and $s_{\textsc{max}}^{\lambda}$ to match the budget $B$.
@@ 3 Cascading as Sequential Routing
In this section, we extend our analysis of the optimal routing strategy to the cascade setting, providing detailed proofs for all the statements in this section in App. B. The solution derived here will be used to develop the optimal strategy for cascade routing in Section 4.
@@
In cascading, an input query is passed through a chain of increasingly larger and more expensive models. The cascade stops once a model’s output meets a certain condition, and that output is returned. We will reinterpret cascading as a sequence of routing problems. To do so, we first define the models over which we need to route, which we refer to as supermodels.
@@
A supermodel $M$ is a sequence of models $(m_{i_{1}},\dots,m_{i_{j}})$ such that running a query through $M$ is equivalent to running it through each of the models in the sequence. The set of all supermodels is denoted as $\mathcal{M}$. By $M_{i:j}$ we denote the supermodel $(m_{i},\dots,m_{j})$.
@@
A cascade occurs as a sequence of decision steps, where at each step, it decides whether to compute the next model in the sequence or terminate. By step $j$, the first $j-1$ models have been computed. At this point, the cascade will run one of the supermodels $M_{1:j-1},\dots,M_{1:k}$. If any of the supermodels $M_{1:j},\dots,M_{1:k}$ is optimal, the cascade proceeds to run the next model. In contrast, if the supermodel $M_{1:j-1}$ is optimal, the cascade halts and returns the output of the last computed model.
Thus, a cascade can be characterized as a sequence of routing strategies that route between the supermodels $M_{1:j-1},\dots,M_{1:k}$. Even though the action associated with supermodels $M_{1:j},\dots,M_{1:k}$ – continuing the cascade – is the same, it is essential to use all of them in the routing strategy. For instance, if $m_{j}$ performs poorly but $m_{j+1}$ performs exceptionally well on a given query, the cascade should continue. Limiting consideration to the supermodels $M_{1:j-1}$ and $M_{1:j}$ alone would therefore result in suboptimal decisions. Formally, we define a cascading strategy as follows:
@@
A cascading strategy $s$ is a sequence of routing strategy $(s^{(1)},\dots,s^{(k)})$ such that $s^{(j)}$ routes between the supermodels $M_{1:j-1},\dots,M_{1:k}$.
@@
We derive the quality and cost estimates associated with supermodel $M_{1:i}$, denoted as $\hat{q}^{(j)}_{1:i}(x)$ and $\hat{c}^{(j)}_{1:i}(x)$, based on the quality and cost estimates of the individual models. Trivially, the cost estimate of the supermodel is equal to the sum of the individual model costs. The quality of a supermodel, however, is governed by the best model within it. Thus, it equals $\mathbb{E}[\max(\hat{q}_{1}(x),\dots,\hat{q}_{i}(x))]$, where the expected value reflects the uncertainty in each quality estimate. This expected value is crucial since ignoring uncertainty would falsely assume that the quality of a supermodel is always equal to the best model within it, even though the best model may return a poor answer, while another returns a good one. To estimate the uncertainties associated with the quality estimates, we compute the variance of $\hat{q}^{(j)}_{i}(x)-\hat{q}^{(k)}_{i}(x)$ over a validation dataset $D$.
While the form of these estimates works well in practice and is intuitively clear, we note that alternative approaches are possible. However, the optimality proof of our cascading strategy does not depend on the specific form of the estimates, and can therefore be adapted to other formulations.
@@
We now leverage the optimal routing strategy from Theorem 1 to determine the optimal cascading strategy. As before, optimality is defined in terms of maximizing the expected output quality while adhering to a given cost budget. However, the budget is now only enforced over the entire cascade, and not over the individual steps. This leads to a slightly different formulation of the optimal cascading strategy:
@@
For a given cost budget $B$, there exist $\lambda_{1},\dots,\lambda_{k}\in\mathbb{R}^{+}$ and a $\gamma\in[0,1]$ such that the optimal cascading strategy $s_{\textsc{opt}}=(s^{(1)}_{\textsc{opt}},\dots,s^{(k)}_{\textsc{opt}})$ is given by the equalities $s^{(j)}_{\textsc{opt}}=\gamma s^{(j),\lambda_{j}}_{\textsc{min}}+(1-\gamma)s^{(j),\lambda_{j}}_{\textsc{max}}$ where $s^{(j),\lambda_{j}}_{\textsc{min}}$ and $s^{(j),\lambda_{j}}_{\textsc{max}}$ are defined as in Theorem 1.
@@
We could not find an analytical solution to this problem. Therefore, we assume $\lambda_{1}=\dots=\lambda_{k}=\lambda$ and apply the binary search technique from Section 2 to determine the optimal $\lambda$. This value is then used as initialization for a hyperparameter optimization tool, Hyperopt [^2] https://github.com/hyperopt/hyperopt, to find the optimal values.
@@
Prior work on cascading has often relied on strong assumptions to simplify the strategy. The most common technique uses a threshold to decide whether to continue the cascade on an input $x$ (Chen et al., 2023; Gupta et al., 2024). Specifically, in step $j$, the cascade continues if $\hat{q}^{(j)}_{j-1}(x)<\tau_{j}$ for some threshold $\tau_{j}\in\mathbb{R}$. Below, we outline the conditions under which this simplified approach is optimal.
@@
Under minor technical assumptions, the thresholding strategy is equivalent to our cascading strategy if and only if the following conditions hold: the cost estimate $\hat{c}^{(j)}_{i}(x)$ is independent of $x$ for all $i,j\in\{1,\dots,k\}$, $\hat{q}^{(j)}_{i}(x)$ is independent of $x$ for all $i\geqslant j$, and the quality estimate $\hat{q}^{(j)}_{1:i}(x)$ is equal to $\hat{q}^{(j)}_{i}(x)$.

### v2

@@ Abstract
The availability of a wide range of large language models (LLMs) embedded in various agentic systems has significantly increased the potential of model selection strategies to improve the cost-performance tradeoff. Existing strategies involve either routing, where a single model is chosen per query, or cascading, which sequentially runs increasingly larger models until a satisfactory answer is found. However, current approaches face three key limitations: they (1) lack formal proofs of optimality, (2) fail to identify the conditions under which these strategies are most effective to improve the cost-performance tradeoff, and (3) are unable to combine both paradigms for further improvements. To address these issues, we first derive a novel optimal strategy for cascading and prove the optimality of an existing routing strategy. Further, we propose cascade routing, a unified framework that integrates routing and cascading into a theoretically optimal strategy. Through our analysis, we identify good quality estimators as the critical factor for the success of model selection paradigms. Finally, in our experiments, we show that cascade routing consistently outperforms the individual approaches by a large margin and we analyze quality estimators to determine when routing and/or cascading are useful paradigms for model selection. [^1] Code available at https://github.com/eth-sri/cascade-routing
LLMs, Routing, Cascading
@@ 1 Introduction
Two primary strategies have been proposed to solve model selection. The first, routing, directs each input query to a specific model from a set of available models (Chen et al., 2022; Liu et al., 2024), as illustrated in Fig. 1(a). This approach is particularly effective when different expert LLMs are needed for diverse tasks, enabling the selection of the most suitable expert for each query. The second strategy, cascading, processes an input query through a sequence of increasingly larger models, stopping when a model produces an answer deemed sufficiently good (Chen et al., 2023; Varshney & Baral, 2022), as illustrated in Fig. 1(b). Cascading is particularly valuable for handling queries of varying difficulty, as it allows simpler queries to be addressed by smaller models while reserving more complex queries for larger models.
@@
Restrictive Conditions
Despite their utility, both routing and cascading impose significant restrictions on the model selection process. In routing, the initial selection of a model is final, preventing any reconsideration after the initial decision. In cascading, each query must sequentially pass through all models in the chain, with no option to skip a model. Therefore, a less restrictive strategy that combines the strengths of both routing and cascading could offer significant performance improvements.
Lack of Deeper Understanding of Routing Algorithms
Further, the conditions under which current routing and cascading strategies are optimal, are not well understood. For routing, an extensive proof is required just to show that current strategies are close to optimal (Chen et al., 2022), while the theoretical analysis of current cascading algorithms is based on restrictive assumptions and does not provide optimality guarantees (Chen et al., 2023; Varshney & Baral, 2022). This lack of theoretical understanding hinders the development of more effective model selection strategies. More importantly, prior work fails to provide insights into the limitations of model selection strategies and cannot identify the conditions under which they are useful in practical scenarios. For instance, it is widely believed that one needs models of various sizes to benefit from cascading (Chen et al., 2023; Gupta et al., 2024; Khalili et al., 2022), but we show that this notion is incorrect.
@@
To address these limitations, we first derive optimal routing and cascading strategies by framing them as linear optimization problems aimed at maximizing output quality while remaining within a given cost budget. For routing, this optimal strategy is close to the one obtained by prior work, while for cascading we derive a new strategy that is provably better than existing approaches. Building on this analysis, we propose a new paradigm called cascade routing, which generalizes both routing and cascading. As illustrated in Fig. 1(c), cascade routing initially routes a query to any available model but keeps rerouting to different models until a model produces an answer of sufficient quality. We prove the optimality of cascade routing and show that it offers significantly more flexibility in processing a query.
Importance of Quality Estimation
Our theoretical analysis enables a more thorough investigation into the factors that influence the effectiveness of model selection strategies. Specifically, we find that accurate estimates of model performance and response quality are most important. For routing, reliable ex-ante quality estimation—the ability to predict whether a model will perform well on a given query—is essential. For cascading, robust post-hoc quality estimation—the ability to evaluate the quality of a model’s response after generation—is critical. Without it, the effectiveness of cascading strategies is severely limited even when models of various sizes are available.
@@ Results
We evaluate cascade routing on a range of tasks, demonstrating that it significantly outperforms both routing and cascading. Notably, cascade routing consistently outperforms other methods, improving performance by up to $8\%$ on the RouterBench benchmark (Hu et al., 2024) and by $14\%$ on the SWE-Bench benchmark (Jimenez et al., 2024). Further, we show that our new cascading strategy outperforms existing cascades in several scenarios by over $10\%$.
Key Contributions
Our main contributions are:
@@
  • We conduct a thorough evaluation, demonstrating that cascade routing consistently outperforms the baselines, highlighting the critical role of quality estimation for the effectiveness of model selection (Section 5).
@@ 2 Routing as Linear Optimization
We derive an optimal routing strategy to select the best model for a given query, providing detailed proofs for all statements in this section in App. A.
Brief Overview
In this section, we begin by defining a routing strategy as a function that maps queries to models. Next, we introduce the notation for quality and cost functions, demonstrating how the optimal routing strategy can be formulated to maximize a linear tradeoff between these two factors. Following this, we explain the process of selecting the best hyperparameters for the strategy. Lastly, we discuss the significance of quality estimation in routing and its impact on the effectiveness of the strategy.
Routing
@@
A routing strategy selects a model by sampling from the distribution $s(x)$ for each query $x$. In prior work, routing strategies were restricted to be deterministic, i.e., $s_{i}(x)\in\{0,1\}$ (Chen et al., 2022; Hu et al., 2024). In contrast, we propose using a more general probabilistic strategy that enables a better solution and an easier theoretical analysis.
@@
However, since these functions are unknown in practice, we need estimators $\hat{q}_{i}(x)$ and $\hat{c}_{i}(x)$ that approximate the output quality and cost of querying model $m_{i}$ on input $x$. Estimators for $q_{i}(x)$ can be created using small classifiers trained to predict model accuracy, as done in prior work (Hu et al., 2024; Shnitzer et al., 2023). $\hat{c}_{i}(x)$ can be estimated by tokenizing the input query and determining the average output length of the model on a query. Then, we can use API-specific costs per token to estimate the cost of running a model on a query. Alternatively, we can also use average execution time as a proxy for cost.
@@
The optimal routing strategy $s_{\textsc{opt}}$ for a given cost budget $B$ is the solution to the optimization problem that maximizes the expected output quality of the selected model while adhering to the budget:
$\max_{s}$ $\mathbb{E}_{x\in\mathcal{X}}\left(\sum\nolimits_{i=1}^{k}s_{i}(x)\hat{q}_{i}(x)\right)$ (1)
s.t. $\mathbb{E}_{x\in\mathcal{X}}\left(\sum\nolimits_{i=1}^{k}s_{i}(x)\hat{c}_{i}(x)\right)\leqslant B.$
We now explain how to solve this optimization problem. For a given query $x$, it can be shown (see App. A) that the optimal routing strategy selects the model maximizing the cost-quality tradeoff $\tau_{i}(x,\lambda)=\hat{q}_{i}(x)-\lambda\hat{c}_{i}(x)$. Here, $\lambda\in\mathbb{R}^{+}$ is a hyperparameter that controls the balance between quality and cost based on the budget $B$.
However, it can occur that several models achieve the same optimal cost-quality tradeoff for a given query. To address this, we define two deterministic strategies $s_{\textsc{min}}^{\lambda}(x)$ and $s_{\textsc{max}}^{\lambda}(x)$, which, respectively, select the cheapest and most expensive model that achieves the optimal tradeoff. The optimal routing strategy $s_{\textsc{opt}}$ is then determined by:
@@
For a cost budget $B$, there exists a $\lambda\in\mathbb{R}^{+}$ and a $\gamma\in[0,1]$ such that the optimal routing strategy $s_{\textsc{opt}}$ equals $\gamma s_{\textsc{min}}^{\lambda}+(1-\gamma)s_{\textsc{max}}^{\lambda}$.
Theorem 1, continued. Furthermore, all routing strategies that have an expected cost that is exactly equal to $B$ and can be written as a convex combination of $s_{\textsc{min}}^{\lambda^{\prime}}$ and $s_{\textsc{max}}^{\lambda^{\prime}}$ for some $\lambda^{\prime}\in\mathbb{R}^{+}$ achieve the same optimal quality.
Since $\gamma$ is often not equal to 00 or $1$, $s_{\textsc{opt}}$ is not necessarily deterministic, i.e., there are queries $x$ such that $s_{\textsc{opt},i}(x)\notin\{0,1\}$ for some index $i$. Therefore, prior work that only considered deterministic routing strategies (Chen et al., 2022; Hu et al., 2024) cannot express the optimal routing strategy and fall back to the near-optimal $s_{\textsc{min}}^{\lambda}$.
Due to the second part of Theorem 1, we only need to find a set of hyperparameters $\lambda$ and $\gamma$ that achieve the cost budget. To determine these parameters, we estimate the cost of a strategy using a validation dataset $D$ that is representative of the query distribution $\mathcal{X}$. We then perform a hyperparameter search to find optimal values of $\lambda$ and $\gamma$. By leveraging several properties of routing strategies (see App. A), one can show that this hyperparameter search can be reduced to a single binary search over $\lambda$, enabling a quick and efficient hyperparameter optimization process.
Ex-Ante Quality Estimation
We explicitly distinguish between the model’s true quality $q_{i}(x)$ and the ex-ante quality estimate $\hat{q}_{i}(x)$. An optimal routing strategy will select a good model only if $q_{i}(x)\approx\hat{q}_{i}(x)$, otherwise the objective in Eq. 1 is not appropriate. Thus, even when routing is suited for the application, the strategy will fail if the quality estimates are inaccurate. This makes quality estimation a critical component of routing strategies. While cost estimation also faces similar challenges, we found that it is less critical and can be approximated more easily.
@@ 3 Cascading as Sequential Routing
We extend our analysis of the optimal routing strategy to a cascade, providing proofs for all statements in App. B.
Brief Overview
In this section, we reinterpret cascading as a sequence of routing problems. We then derive the optimal cascading strategy and explain how to determine the optimal hyperparameters by solving an additional optimization problem. Furthermore, we show how our approach improves upon the cascading strategies used in prior work. Finally, we examine the impact of post-hoc quality estimates on the effectiveness of cascading strategies.
@@
In cascading, an input query is processed sequentially through a chain of models, typically arranged in order of increasing size or cost. The cascade stops once a model’s output meets a certain condition, and that output is returned. We will reinterpret cascading as a sequence of routing problems. To do so, we first define the models over which we need to route, which we refer to as supermodels.
@@
A supermodel $M$ is a sequence of models $(m_{i_{1}},\dots,m_{i_{j}})$ such that running a query through $M$ is equivalent to running it through each of the models in the sequence. $\mathcal{M}$ denotes the set of all supermodels and by $M_{i:j}$ we denote the supermodel $(m_{i},\dots,m_{j})$.
@@
Cascading as Sequential Routing
Running a cascade on a sample $x$ occurs in a sequence of steps, where at each step, the cascade determines whether to run the next model in the sequence or terminate. By step $j$, we have obtained outputs from the first $j-1$ models. To decide whether to continue and run $m_{j}$, we need to determine, in expectation, how well the supermodels $M_{1:j-1},\dots M_{1:k}$ will perform on the sample $x$. Once again, this performance is measured as having the highest expected output quality within a certain cost budget. If $M_{1:j-1}$ offers the best performance, we terminate the cascade and return its output, i.e., the output of $m_{j-1}$. Otherwise, if any of $M_{1:j},\dots,M_{1:k}$ has better performance, we continue the cascade and run $m_{j}$. Therefore, at step $j$, the cascade is equivalent to a routing strategy that selects the best supermodel from $M_{1:j-1},\dots,M_{1:k}$. Thus, a cascade can be formally defined as follows:
@@
A cascading strategy $s$ is a sequence of routing strategies $(s^{(1)},\dots,s^{(k)})$ such that $s^{(j)}$ routes between the supermodels $M_{1:j-1},\dots,M_{1:k}$.
Notably, a model $m_{j}$ might perform poorly while $m_{j+1}$ performs exceptionally well on a given query. In such cases, the performance of $M_{1:j}$ will be worse than the performance of $M_{1:j-1}$, but $M_{1:j+1}$ could still provide a better outcome. Therefore, it is crucial to consider all supermodels $M_{1:j},\dots,M_{1:k}$ at step $j$ rather than making decisions solely based on immediate performance.
@@
We derive the quality and cost estimates associated with supermodel $M_{1:i}$, denoted as $\hat{q}^{(j)}_{1:i}(x)$ and $\hat{c}^{(j)}_{1:i}(x)$, based on the quality and cost estimates of the individual models. Trivially, the cost of the supermodel is equal to the sum of the individual model costs. The quality of a supermodel, however, is governed by the best model within it. Thus, it equals $\mathbb{E}_{\hat{q}_{i}}[\max(\hat{q}_{1}(x),\dots,\hat{q}_{i}(x))]$, where the expected value reflects the uncertainty in each quality estimate. Specifically, each quality estimate $\hat{q}_{i}(x)$ is modeled as a random variable estimating the true quality $q_{i}(x)$. This is crucial since ignoring uncertainty would falsely assume that the quality of a supermodel is always equal to the best model within it, even though the best model may return a poor answer, while another returns a good one. To estimate the uncertainties associated with the estimates, we compute the variance of $\hat{q}^{(j)}_{i}(x)-\hat{q}^{(k)}_{i}(x)$ over a validation dataset.
@@
We now leverage the optimal routing strategy from Theorem 1 to determine the optimal cascading strategy. As before, optimality is defined in terms of maximizing the expected output quality while adhering to a given cost budget. However, the budget is now only enforced over the entire cascade, and not over individual steps. This leads to a slightly different formulation:
@@
For a given cost budget $B$, there exist $\lambda_{1},\dots,\lambda_{k}\in\mathbb{R}^{+}$ and a $\gamma\in[0,1]$ such that the optimal cascading strategy $s_{\textsc{opt}}=(s^{(1)}_{\textsc{opt}},\dots,s^{(k)}_{\textsc{opt}})$ is given by $s^{(j)}_{\textsc{opt}}=\gamma s^{(j),\lambda_{j}}_{\textsc{min}}+(1-\gamma)s^{(j),\lambda_{j}}_{\textsc{max}}$ where $s^{(j),\lambda_{j}}_{\textsc{min}}$ and $s^{(j),\lambda_{j}}_{\textsc{max}}$ are defined as in Theorem 1.
@@
To solve this, we first assume $\lambda_{1}=\dots=\lambda_{k}=\lambda$ and apply the binary search technique from Section 2 to determine $\lambda$. This is then used as initialization for a hyperparameter optimization tool, Hyperopt [^2] https://github.com/hyperopt/hyperopt, to find the optimal values.
@@
Prior work on cascading has often relied on strong assumptions to simplify the strategy. The most common technique uses a threshold to decide whether to continue the cascade on an input $x$ (Chen et al., 2023; Gupta et al., 2024). Specifically, in step $j$, the cascade continues if $\hat{q}^{(j)}_{j-1}(x)<\tau_{j}$ for some $\tau_{j}\in\mathbb{R}$. Below, we outline the conditions under which this simplified approach is optimal.
@@
Under minor technical assumptions, the thresholding strategy is equivalent to our cascading strategy if and only if the following conditions hold: $\hat{c}^{(j)}_{i}(x)$ is independent of $x$ for all $i,j\in\{1,\dots,k\}$, $\hat{q}^{(j)}_{i}(x)$ is independent of $x$ for all $i\geqslant j$, and $\hat{q}^{(j)}_{1:i}(x)$ is equal to $\hat{q}^{(j)}_{i}(x)$.
Post-Hoc Quality Estimation
Once again, this theoretical framework highlights the importance of quality estimation, with a shift in focus from ex-ante quality estimation to post-hoc quality estimation, which now plays a critical role. Cascading approaches are only advantageous when the post-hoc quality estimate provides significantly better information than the ex-ante estimate. If this improvement is minimal, it would be more effective to directly route queries to the most suitable model, bypassing the cascading process. While the post-hoc estimate is essential for refining decisions, the ex-ante quality estimate remains valuable in determining whether future models can potentially deliver better performance. Only in the threshold cascade strategy, where the ex-ante estimate is fixed, does it become irrelevant. By contrast, our approach improves upon the threshold cascade by incorporating both ex-ante and post-hoc quality estimates, thereby enabling more informed decision-making.